Write a system of two equations in two unknowns for each problem Solve the systems by the method of your choice.
At Gwen’s garage sale, all books were one price, and all magazines were another price. Harriet bought four books and five magazines for $1.45, and June bought two books and five magazines for $1.25. What was the price of a book and what was the price of a magazine?

Answer by checkley71(8403)

145=4B+5M & 125=2B+5M NOW MULTIPLY THE SECOND EQUATION BY -1 WE GET -125=-2B-5M
NOW WE ADD THESE TWO EQUATIONS
145=4B+5M
-125=-2B-5M
---------------------
20=2B OR B=20/2 OR B=10 CENTS THEN 145=4*10+5M OR 145=40+5M OR 5M=145-40 OR
5M=105 OR M=105/5 OR 21 CENTS FOR EACH MAGAZINE.